Given υ(t) = 25 + 18t, where υ is in m/s and t is in s, use calculus to determine the total displacement  

from t1 = 1.5 s to t2 = 3.1 s. [ANS = 106 m]


1
Expert's answer
2021-11-07T19:30:05-0500
"d=\int^{3.1}_{1.5}(25+18t)\text dt=25t+9t^2\bigg|^{3.1}_{1.5}=106.24\text{ m}."